OpenStudy (anonymous):

substitution x-y=7 and 3x+2y=6 PLEASE SHOW THE WORK thank you and i will best response your answer :)

OpenStudy (anonymous):

x-y=7. change it to y=x-7. input that as y. 3x+2(x-7)=6. that would simplify to 3x+2x-14=6. then simpiply more to 5x=20. then x=4.

OpenStudy (anonymous):

thank you thanks you thank you :)